### v1

# A mix of long-duration hydrogen and thermal storage enables large-scale electrified heating in a renewable European energy system

Abstract
Hydrogen-based long-duration electricity storage (LDES) is a key component of renewable energy systems to deal with seasonality and prolonged periods of low wind and solar energy availability. In this paper, we investigate how electrified heating with heat pumps impacts LDES requirements in a fully renewable European energy system, and which role thermal storage can play. Using a large weather dataset of 78 weather years, we find that electrified heating significantly increases LDES needs, as optimal average energy capacities more than quadruple across all weather years compared to a scenario without electrified heating. We attribute 75% of this increase to a leverage effect, as additional electric load amplifies storage needs during times of low renewable availability. The remaining 25% are the result of a compound effect, where exceptional cold spells coincide with periods of renewable scarcity. Furthermore, heat pumps increase the variance in optimal storage capacities between weather years substantially because of demand-side weather variability. Long-duration thermal storage attached to district heating networks can reduce LDES needs by on average 36%. To support and safeguard wide-spread heating electrification, policymakers should expedite the creation of adequate regulatory frameworks for both long-duration storage types to de-risk investments in light of high weather variability.
journal: arXiv
[…]
references.bib
LDES
long-duration electricity storage
LDTS
long-duration thermal storage
TWh
terawatt-hours
DH
district heating
COP
coefficient of performance
MCRL
maximum cumulative residual load
VRE
variable renewable energy
1 Introduction
The global effort to mitigate climate change relies crucially on the decarbonization of the energy system [intergovernmental_panel_on_climate_change_ipcc_technical_2023, international_energy_agency_net_2024]. For many countries, especially in Europe, studies suggest that a power system largely based on variable wind and solar power is feasible and cost-effective to decarbonize the energy system [brown_response_2018, hansen_status_2019, luderer_impact_2022, breyer_history_2022]. Planning weather-resilient variable renewable energy (VRE) systems requires considering weather variations over long timescales and extreme events [pfenninger_dealing_2017, van_der_wiel_influence_2019, bloomfield_quantifying_2016, craig_overcoming_2022]. A growing strand of the literature analyzes the impact of weather variability on optimal renewable system designs [gotske_designing_2024, ruggles_planning_2024, grochowicz_intersecting_2023, goke_stabilized_2024], investigates periods of extremely low renewable generation, also referred to as Dunkelflaute [kittel_measuring_2024, kittel_coping_2025], determines flexibility needs to deal with variability and extreme events [lund_review_2015, alizadeh_flexibility_2016, brown_synergies_2018, dowling_role_2020, ruhnau_storage_2022], and seeks to overcome the computational challenges associated with including decades of weather data in capacity expansion models [goke_stabilized_2024, jacob_future_2023].
Energy storage technologies are central for bridging periods with insufficient supply of solar or wind energy [schill_electricity_2020, levin_energy_2023]. To provide energy during prolonged Dunkelflaute events, long-duration electricity storage (LDES) has emerged as a key enabler [levin_energy_2023, dowling_role_2020, dowling_long-duration_2021, staadecker_value_2024, jenkins_long-duration_2021]. While a range of LDES technologies is potentially available, the key metric for their competitiveness is the cost of energy capacity [sepulveda_design_2021]. Despite geographic restrictions that may put a strain on transmission infrastructure [brown_ultra-long-duration_2023], the debate has been centered around hydrogen cavern storage [dowling_long-duration_2021]. Model results on optimal LDES capacities vary substantially on the selected weather data [kittel_coping_2025, ruhnau_storage_2022], geographic scope, considered sectors, and other model specifications.
The direct or indirect electrification of sectors such as transport, industry, and heating is crucial for the complete decarbonization of economies around the world [international_energy_agency_net_2024]. Yet, most studies at the nexus of LDES and weather variability focus only on the power sector and neglect the impact of weather variations on the demand side of a sector-coupled renewable system [gotske_designing_2024]. Accounting for 16% percent of total greenhouse gas emissions in the European Union in 2024 [rozsai_m_jrc-idees-2021_2024], the heating sector takes center stage in the continent’s decarbonization efforts [commission_renewable_2023]. Previous studies have shown that serving a substantial part of residential and commercial heat demand with heat pumps can have challenging consequences for the power system [bloess_power--heat_2018, zeyen_mitigating_2021, roth_power_2023, roth_power_2024]. Furthermore, there is a substantial degree of interannual variation in temperatures, and correspondingly in heat demand and temperature-dependent heat pump efficiencies. There is evidence that cold spells can coincide with times of renewable supply shortages [van_der_wiel_influence_2019, mockert_meteorological_2023] exacerbating system stress.
[…]
LDTS is a potential alternative to LDES to deal with heating-related system imbalances. Thermal energy storage has been shown to improve the utilization of renewables and to reduce heat prices by smoothing heat demand peaks in Danish case studies [sifnaios_impact_2023, christensen_role_2024]. Another Danish study finds it to be a no-regret option for flexibility in district energy systems with electrolyzers [ostergaard_optimal_2023]. Short-duration thermal storage can complement batteries [jacob_future_2023, roth_power_2023]. Zeyen et al. conclude that thermal storage can mitigate heating demand peaks in a sector-coupled European energy system, especially in combination with building renovations [zeyen_mitigating_2021]. A Norwegian case study discusses alternative configurations of long-duration thermal storage (LDTS) in local district heating systems with renewables [kauko_flexibility_2024]. A direct comparison of hydrogen and thermal storage for the Swedish system concludes that hydrogen storage is better suited to integrate wind power, but less effective in reducing fuel consumption [sundarrajan_harnessing_2025]. In a large sector-coupled European system model for the weather year 2015, Victoria et al. find that electrifying the heating sector comes with large LDTS capacities, but these do not replace hydrogen storage capacities [victoria_role_2019]. In conclusion, the the previous literature on heat electrification and LDTS is largely focused on local energy systems and usually considers single or few weather years only.
We contribute to the literature with an integrated analysis of long-duration electricity and thermal storage needs in a fully renewable European energy system with large-scale electrified heating. We do so at an extensive geographic and temporal scope, using a sector-coupled capacity expansion model spanning 28 European countries at an hourly resolution and using a comprehensive dataset of 78 weather years. The latter features weather-dependent time series of wind and solar capacity factors, hydro inflows, heat demand and heat pump efficiencies. We exploit the variation between weather years to isolate the drivers of increased LDES needs, differentiating between a leverage and a compound effect. We further quantify the potential role of LDTS in providing long-duration flexibility and discuss its potential to substitute or complement LDES, while most of the existing literature has focused on short-duration thermal energy storage. Based on our findings, we discuss implications for potential storage investors and policymakers.
2 Results
In the following, we focus primarily on the link between electrification of residential and commercial heat and LDES needs. The Supplemental Information Section SI.4 provides several additional results. We summarize all considered scenarios in Table 1 and provide additional details in Section 4.3.
Table 1: Scenarios
Scenario Heat demand LDTS Description
No Heat - - Historical electricity demand corrected for electric heating
Decent Year-specific - 80% of heat demand supplied by decentralized heat pumps
Decent - Mean Long-run mean - Like Decent but with mean heat demand profile
District & Decent Year-specific Yes 32% of heat demand supplied by district heating enabling LDTS
2.1 Electrified heat demand leads to significantly higher long-duration electricity storage requirements
To identify how much additional LDES energy capacity is needed due to the electrification of heat, we define two sets of scenario runs. In the first set (No Heat), the model jointly optimizes capacity and dispatch decisions in a fully renewable system without any electrified heat. The resulting aggregate LDES energy capacities for all 28 countries vary by weather year (gray dots in Figure 1). Optimal aggregate capacities across Europe range between 23  terawatt-hours (TWh) in 1988/89 to 80 TWh in 1958/59. The average LDES capacity across all weather years is 37 TWh.
In the second set of runs (Decent), we assume that 80% (see Section 4.3) of all residential and commercial heat demand (space and hot water) is electrified, using decentralized heat pumps with small buffer thermal storage. This increases the mean aggregate LDES requirement to 168 TWh, corresponding to an increase of 131 TWh (273%) on average (represented by the dark-red dots in Figure 1). In 1962/63, the year with the highest LDES requirement in the set of runs with electrified heat, the LDES capacity increases tenfold from 40 TWh to 400 TWh.
In Figure 1, the weather years are ordered by LDES requirement in the Decent scenario. Notably, the No Heat scenario does not reflect the same ordering. Weather years that have high LDES requirements in the scenario without electrified heat do not necessarily rank high when considering electrified heat (cf. 1989/90) and vice versa (cf. 1941/42). Additionally, the distribution of optimal LDES capacities widens substantially. In the No Heat scenario, the standard deviation of total LDES energy capacities across all weather years is 10 TWh. In the scenario Decent, the standard deviation increases to 66 TWh.
Figure 1: Long-duration electricity storage requirements in scenarios with different electrified heat demand levels
Optimal long-duration electricity storage capacities by weather year. Grey dots represent optimal capacities in the No Heat scenario. Dark-red dots represent optimal capacities in the Decent scenario. Light-red crosses represent optimal capacities in the Decent - Mean scenario, in which weather-year specific heat demand is replaced by long-run mean heat demand.
2.2 Heat demand drives storage needs through ‘leverage’ and ‘compound’ effects
In a third set of scenario runs (Decent - Mean), we replace the time series of heat demand and heat pump coefficients of performance (COP) in a given weather year by their long-run hourly averages across all weather years in the sample. These curves (see Figure SI.11) isolate the inherent seasonality of electricity demand for heating and abstract from cold spells or unusually warm periods. The results of these model runs (light red crosses in Figure 1) highlight the drivers of increasing LDES requirements. Heat demand is seasonal and electrified heating naturally adds most load in winter months. As extreme system states of VRE scarcity mostly occur during winter months [kittel_quantifying_2024, kittel_coping_2025, gotske_designing_2024], the additional load amplifies the energy deficit, triggering additional LDES requirements by functioning as a lever. Across the sample of weather years, this leverage effect accounts for 98 TWh or 75% of the total LDES increase on average.
For some years, the Decent - Mean scenario results in even higher LDES requirements (e.g., 1973/74) in which unusually mild conditions offset some of this leverage effect. Yet, in other years the year-specific heat demand leads to far higher LDES capacities, as VRE scarcities coincide with cold spells in a compound effect. While this effect only explains 25% on average, cold spells have a considerable impact in some of the most extreme weather years.
Figure 2: LDES trajectories with and without electrified heat in selected weather years
LDES trajectories shown against daily deviations in heat demand from long-run mean (bright red - mid) and 24-hour rolling renewable availability (green - low). Gray area represents storage trajectory in No Heat. The darker red area shows the additional storage level in Decent - Mean, and the lighter red addition represents the incremental storage level from the specific weather year’s heat demand (Decent).
We illustrate both effects for four weather years with high additional LDES requirements in Figure 2. In the deterministic setting addressed here, the maximum required storage level over the course of a weather year determines the optimal LDES energy capacity. The trajectories therefore provide an insight into the critical periods that determine optimal LDES capacities. The gray areas correspond to the storage levels observed in the scenario without electrified heat (No Heat). The darker red area reflects additional storage levels required to meet the mean heat demand profile (Decent - Mean). The lighter red areas show the storage trajectories arising from the weather year’s specific heat demand profile (Decent). In the lower part of the figure, we plot the absolute daily deviations from long-run mean profiles of both heat demand (in dark-red) and renewable generation. The heat deviations serve as a proxy for cold spells. For renewable generation, we take the mean generation portfolio of the Decent scenario as a basis. The deviations are averaged in a 24-hour rolling window and are indicative of renewable energy scarcity, where the area under the zero line can be interpreted as an energy deficit.
We observe the most extreme LDES capacity increase in 1962/63 (left). In the No Heat scenario, the storage level traces four episodes of relative VRE scarcity beginning in October 1962. The leverage effect amplifies all four episodes, particularly so in January when seasonal heat demand reaches its peak. By some definitions, the winter of 1962/63 was the coldest European winter on record since 1739/40 [greatbatch_tropical_2015] with a long, incessant cold spell from December to March encapsulating all renewable scarcity episodes and compounding to a continuous discharge event of close to 334 TWh over the same period.
The winter of 1996/97 (center left), often identified as an extreme weather year for renewable energy systems due to an episode of extremely low wind capacity factors [kittel_quantifying_2024, grochowicz_using_2024], was relatively mild, but included a brief cold snap in January. Featuring the second-highest storage requirements in the No Heat scenario, it is a dominating leverage effect that adds LDES capacity. The cold spell only adds about 70 TWh.
The weather year of 1958/59 (center right) results in the highest LDES capacity in the No Heat scenario due to a severe episode of energy scarcity between November and early January, leading to a discharge event of 79 TWh over the course of 28 days. Rather mild temperatures induce a LDES increase almost entirely through the leverage effect, where the discharge event increases in size to 223 TWh end extends to 106 days.
Lastly, the season of 1941/42 (right) exhibits one of the lowest LDES requirements across all years in the sample when we do not include electrified heat. However, a moderate but prolonged VRE scarcity period combined with a heavy cold spell in the first quarter of 1942 results in the fifth-highest LDES requirements.
We note from these illustrative instances that different combinations of renewable scarcity events and cold spells can lead to significant LDES capacity requirements.
2.3 Long-duration thermal storage in district heating networks can mitigate LDES requirements
[…]
In contrast to a hydrogen cavern storage, a LDTS has relatively high standing losses from heat loss [sifnaios_heat_2023]. In the base specification, we assume an efficiency equivalent to cumulative losses of 39% within 90 days of storage in energetic terms [zeyen_mitigating_2021]. Hydrogen-based electricity storage, which includes electrolysis, cavern storage, and reconversion to electricity with gas turbines, has a relatively low roundtrip efficiency of around 30%, but is independent of the storage duration as standing losses are negligible. Even under these conservative estimates for LDTS efficiency, a LDTS system is more efficient in terms of electrical energy input per unit of thermal energy demand served for storage duration of up to ca. 180 days - longer than the seasonal heating cycle.
Our modeling results confirm such back-of-the-envelope calculations. The introduction of LDTS reduces the additional LDES requirements from electrified heat by 60 TWh (36%) on average (Figure 3, left panel). In the extreme weather year 1962/63 the introduction of co-optimized LDTS capacities in DH networks leads to a reduction of 155 TWh (37%). The Europe-wide average LDTS capacity over all years is 143 TWh, reaching almost 323 TWh in the maximum (Figure 3, center panel).
The right panel of Figure 3 depicts the aggregate daily electricity sector balance for the year 1962/63 under Decent (upper) and District & Decent (lower). Evidently, hydrogen turbine dispatch is significantly reduced in the District & Decent scenario as heating-network-connected LDTS smooths and replaces electricity demand for heating. This effect is depicted by the reduced red area (electricity demand for heat pumps) in the figure.
Total electricity generation from hydrogen turbines decreases from 88 to 53 TWh on average (151 to 78 TWh for 1962/63). Peak generation falls from 65 to 34 GW (237 to 130 GW). Total electricity demand for heating increases slightly from 793 to 794 TWh on average (1038 to 1057 TWh) due to storage losses, but peak electricity demand from the heating sector falls from 351 to 280 GW (564 to 448 GW). These findings hold for almost all weather years.
Figure 3: LDES requirements without and with district heating thermal storage
left: distribution of LDES capacities across weather years under Decent and District & Decent scenarios; center: distribution of LDTS capacities in District & Decent scenario across weather years (center); right: 1962/63 daily electricity sector balance for Decent scenario (upper) and District & Decent scenario (lower).
The introduction of LDTS only slightly reduces generation capacity investments in solar PV (-5% on average, -4.5% for 1962/63) and wind (-1% on average, -2% for 1962/63), see Figure SI.1.
2.4 Central heat storage decentralizes overall long-duration storage infrastructure
Potentials for cavern storage capacities are unevenly distributed across Europe (Figure SI.10, left panel). Where there is potential in the form of rock salt formations in a country, the potential far exceeds any conceivably necessary storage requirement. Any of those countries could technically host all the cavern storage required for the entire continent. Yet, in the Decent scenario around 43% (72 out of 168 TWh) of the total cavern storage capacity is located in Germany on average, compare Figure SI.3. Notably, cavern storage is much more evenly distributed between countries in the No Heat scenario (Figure SI.2, left panel). Hydrogen gas turbines for reconversion are also concentrated in Germany in Decent, while electrolysis capacity is distributed more evenly, across countries on average.
Figure 4 shows the change in the geographical long-duration storage distribution across Europe when district heating networks with LDTS are introduced for the example of 1962/63. Heating demands and district heating potentials vary between countries such that a meaningful comparison of LDTS deployment requires a relative basis. We express LDTS as a share of total demand in district heating networks. LDTS uptake is particularly high in Central Eastern Europe where there is very little LDES capacity. The reduction in LDES capacity is disproportionally high in Germany. Consequently, the share of European LDES capacity in Germany falls to 35% (38 out of 108 TWh) in District & Decent. This masks two effects. For one, Germany has a relatively high district heating potential compared to other cavern storage locations, which means that it can substitute some of its on flexibility requirements locally. For two, German LDES provides flexibility in other countries, particularly in Central Eastern Europe and high LDTS adoption in these countries also reduces German LDES requirements.
Figure 4: Geographical impact of adding long-duration thermal storage in 1962/63
(a) Share of district heating heat demand covered by LDTS; (b) Change in LDES capacity when introducing LDTS; (c) Change in PEM electrolysis when introducing LDTS; (d) Change in H2 turbine capacity when introducing LDTS.
2.5 Sensitivities
We investigate a range of sensitivities for the results. For computational tractability, we limit these runs to the four years discussed above.
We test alternative specifications of LDTS efficiencies and investment costs. For each of the four weather years, we run 42 alternative settings, varying the LDTS standing losses on the x-axis and the investment costs on the y-axis (Figure 5). The base configuration used in all prior scenarios above is in the corner in the bottom-left. In the most favorable parameter configurations LDTS reduces LDES needs by another 19-35% (upper panel). In the case of 1941/42 this would correspond to an LDES requirement of 111 TWh when LDTS is free and efficiency is at 99%. As per the lower panel of Figure 5, such additional reductions are accompanied by large increases in optimal LDTS capacities. For 1941/42, the LDTS capacity almost triples between the base case and the best parameter combination. Based on the conclusion above that even in the base specification heat losses do not match the roundtrip losses of LDES for relevant storage durations, it is not surprising that additional substitution from efficiency gains beyond a threshold is limited. In the top line of each tile LDTS can be built for free. In these cases, the substitution potential with respect to LDES is limited by the country-level DH potentials.
Figure 5: LDTS sensitivity results
Sensitivity of LDTS parameters. The x-axis varies heat losses from the base case to almost no losses in energetic terms, expressed as energy remaining after 90 days in storage. The y-axis varies the assumed investment costs of LDTS by multiplying them with the labeled scalers.
Next, we explore an extreme counterfactual case in which all electrified heating would be supplied through heating networks (Table 2, last row), which makes LDTS available for all electrified heating demand. In such a setting, LDES requirements would be significantly reduced, but still remain above the No Heat scenario. This reflects the fact that LDTS smooths heating electricity demand peaks by shifting part of the load to months outside the heating season, but does not eliminate all additional electricity demand in winter. This still leaves a leverage effect (compare right panel of Figure 3), as described in Section 2.2.
Table 2: Sensitivities
Aggregate European LDES requirements in TWh for selected sensitivities. ’Base case’ refers to the scenarios used in the remaining analysis. ’No H2 network’ assumes that there is no hydrogen network. ’Islands’ assumes that there is no interconnection between countries at all, ’Nuclear’ assumes fixed nuclear generation capacities of 96.9 GW [entso-e_tyndp_2025]. ’All district heating’ assumes that all electrified heating can be supplied in heating networks with LDTS.
1962/1963 1996/1997 1958/1959 1941/1942
No Heat Base case 40 61 80 27
Decent Base case 400 341 238 278
No H2 network 433 387 313 315
Islands 407 363 318 327
Nuclear 359 282 199 247
District & Decent Base case 252 237 186 171
No H2 network 301 293 268 221
Islands 315 300 274 243
Nuclear 219 180 152 144
All district heating 137 128 136 98
We also test the impact of different interconnection settings. First, we assume that there is no hydrogen network, see rows 3 and 7 of Table 2. Here, the increase in LDES energy capacities is more pronounced in the District & Decent scenario (20% - 44%) than in the Decent scenario (8% - 32%). In the District & Decent base case, hydrogen imports with long-term contracts are actually slightly higher than in the Decent base case, which is likely to replace some LDES capacity otherwise used for serving flat industrial hydrogen demand. In absence of hydrogen import capacities, this option no longer exists.
In the scenario with neither hydrogen nor electricity interconnection capacity (’Islands’, rows 4 and 8), LDES capacities increase only marginally compared to the ’No H2 network’ case and even decrease in the case of 1996/97 in the Decent scenario, indicating that the system value of LDES depends on interconnection. Without interconnection, building larger generation portfolios is cheaper than building expensive hydrogen tank storage. Only about 10 TWh of tank storage are being built across Europe in the ’Islands’ case of the Decent scenario.
Lastly, we test the effect of introducing around 97 GW of nuclear capacity, of which the majority is located in France (58 GW). In either scenario, it has a limited effect on LDES capacities (-14% and -18% on average). The reason is likely to be twofold. For one, nearly two thirds of the capacity being in France, limited interconnection capacities may prevent spatial balancing during extreme system episodes. By contrast, LDES leaves the option to transport both electricity and hydrogen. For two, as Section 2.3 shows, peak supply by hydrogen turbines in winter periods clearly exceeds the aggregate nuclear capacity in extreme years, and hence the introduced nuclear capacities can only supply a limited part of the positive residual load.
3 Discussion and conclusion
3.1 Summary
In this analysis, we isolate the effect of the electrification of heating with heat pumps on optimal long-duration storage needs in a fully renewable European energy system using a large sample of 78 weather years. We find that heat pumps significantly increase long-duration electricity storage (LDES) requirements, mainly because additional electricity demand has to be served in periods of renewable energy scarcity, which typically occur in the heating season during winter [mockert_meteorological_2023, kittel_quantifying_2024]. This leverage effect accounts for 75% of the average additional long-duration electricity storage (LDES) requirements. The remaining 25% can be explained by a compound effect, describing the coincidence of the aforementioned periods of renewable energy scarcity and cold spells, which are periods of exceptionally high heating demand. This leads to considerable positive residual load events that have to be covered by LDES. Similar compound effects are found in previous analyses of extreme system states [gotske_designing_2024, mockert_meteorological_2023].
Previous literature on LDES has noted the sensitivity of optimal LDES capacities to interannual weather variability [dowling_role_2020, ruhnau_storage_2022, kittel_coping_2025]. Yet, these studies have only considered supply-side variability across weather years. Our results suggest varying load, driven by temperature-dependent heating demand, has a much greater impact on optimal LDES deployment. The impact of demand-side weather variability further highlights the necessity to include large datasets in energy system models when planning weather-resilient renewable energy systems.
The use of long-duration thermal storage (LDTS) can effectively reduce the need for LDES but is likely confined to district heating networks. Using an optimistic estimate of the potential for future district heating networks, we find that even a conservatively parameterized LDTS can substitute up to 36% of the LDES requirement in a least-cost system.
While the long-duration storage infrastructure is concentrated in Germany in a scenario with only decentralized heat pumps, the introduction of LDTS leads to a more even geographic distribution of storage capacities. With its limited hydrogen cavern potential, LDTS in Central Eastern Europe can relieve some cavern storage needs in Germany.
Our findings are fairly robust to a range of sensitivities. Cost reductions and efficiency gains for LDTS would further mitigate LDES requirements. Notably, policy-relevant capacities of nuclear power in Europe mitigate LDES and LDTS needs only to a limited extent. Likewise, alternative assumptions on interconnection capacities have only minor impacts on results. For very optimistic alternative assumptions on the costs, efficiency and potentials of LDTS, we find that thermal storage mitigates hydrogen storage needs to a much larger extent than under baseline assumptions.
3.2 Limitations
The large geographic and temporal scope of our analysis comes at the expense of simplifications in the representation of several model components. Specifically, we only use one type of air-sourced heat pump for decentralized heating and do not include any efficiency improvements in the building stock. Similarly, we aggregate all district heating networks within a country and only consider one large-scale heat pump to supply these, abstracting from other potential heat sources. We further use a simplified representation of the thermodynamics of large-scale pit thermal storage, which neglects that some electricity might be required at the time of LDTS discharging to elevate the temperature level again, which could reduce the cost advantage of thermal storage against hydrogen storage. While some of these assumptions (e.g., only air-sourced heat pumps, no other heat sources) lead to a higher electricity demand and higher LDES needs than in reality expected, others might lead to an underestimation (simplified representation of the thermodynamics). Finally, the perfect foresight assumption for a given weather year leads to ideal but unrealistic storage operations and is likely to underestimate the relative value of LDES.
3.3 Conclusions & outlook
In our most challenging setting with 80% electrified heat demand, in the absence of long-duration thermal storage, and considering an extremely cold and long winter such as 1962/63, we find a European long-duration electricity storage (LDES) capacity requirement of 400 TWh. While this capacity is very large compared to currently installed electricity storage, it appears technically feasible considering geological potentials for caverns in Europe [caglayan_technical_2020]. For context, the currently existing natural gas storage capacity in the EU is around 1,100 TWh [agency_for_the_cooperation_of_energy_regulators_report_2022]. However, only a part of the existing gas caverns are suitable for hydrogen storage due to the more demanding geological requirements [oni_underground_2025] and the lower volumetric density of hydrogen compared to natural gas reduces the energy storage capacity. Presently, bar a few test sites, hydrogen cavern storage is a nascent technology.
Long permitting processes and underdeveloped supply chains, as well as long-term uncertainty about hydrogen prices and quantities are likely to increase project risks [steinbach_future_2024]. Similarly, neither electrolysis nor hydrogen gas turbines have been deployed at scale so far [odenweller_green_2025]. Moreover, uncertainties around the ramp-up of heat electrification and hydrogen transmission infrastructure complicate the economics of highly centralized long-duration electricity storage (LDES). In this context, LDTS could provide a promising complementary technology, as it does not rely on particular geological conditions and can effectively deal with long-duration flexibility requirements in the heating sector. Yet, it is likely to be confined to heating networks and our results suggest that even if they were substantially expanded, a sizable residual long-duration flexibility need from the heating sector would remain.
Our results indicate that LDES in the form of cavern storage would be concentrated in a few countries, particularly in Germany, and thus other countries without salt cavern potentials to store hydrogen would be dependent on imports and transmission infrastructure to cover their load and heat demand. As this could raise regulatory and geopolitical questions, LDTS could for some countries increase their energy autonomy and decrease dependence on energy transport infrastructure.
Accordingly, LDES and LDTS should not be considered as substitutes, but rather complement each other, and policymakers should work to enable the mass deployment of both storage types. In particular, policymakers and regulators should expedite the process of designing adequate regulatory frameworks that enable a timely capacity build-out, as a lack of long-duration flexibility would jeopardize the electrification of the heating sector.
From an investor’s perspective, the increased variance of optimal LDES capacities between weather years when considering electrified heating raises additional questions around predictable cost recovery without other economic incentives. This clearly calls for de-risking such investments. Compared to hydrogen storage, it might be easier to manage such risks for thermal storage, as the latter is likely to be developed and operated by vertically integrated, cost-regulated regional monopolists who should be better able to manage quantity and price risks.
Our results show that not only supply-side, but also demand-side weather dependency has a sizable effect on renewable energy systems. Energy modelers and system planners should take this into account when investigating weather-resilient renewable energy systems. To fully capture not only the leverage but also the compound effect, as many weather years as possible should be considered in respective analyses.
As for future research, we consider a more granular representation of the heating sector in combination with a finer geographical resolution a promising avenue. In particular, including alternative heat supply sources such as solar thermal or waste heat from hydrogen turbines or other sources could impact the optimal shares of centralized LDES and local LDTS. Building renovation and the trade-off with storage [zeyen_mitigating_2021] should be appraised in light of interannual weather variability.
Finally, the use of deterministic models omits a potentially important advantage of LDES over LDTS in settings without perfect foresight. LDES can serve final energy demand in many different sectors depending on the realized system state thanks to the usefulness and convertibility of hydrogen. In contrast, LDTS is less useful to the energy system as it can only serve heat demand. Future research, using stochastic modelling (such as in [schmidt_long-duration_2025]) could be a starting point to better understand the value of different types of storage, the interaction between them, and their operation in a non-deterministic setting. However, such approaches are likely to be feasible only for less extensive geographical and temporal scopes than modeled here.
4 Methods
This paper uses a sector-coupled capacity expansion model of Europe across a sample of 78 weather years (1940/41 - 2017/18) at hourly resolution in a scenario analysis to isolate the effect of electrified heating on long-duration storage requirements.
4.1 Sector-coupled energy system model
The energy system model used is an extended version of the open-source power system model DIETER [zerrahn_long-run_2017], written in Julia. The model has been used in various studies to investigate flexibility in sector-coupled VRE-based energy systems, e.g. [zerrahn_economics_2018, schill_electricity_2020, roth_geographical_2023, roth_power_2024, gueret_impacts_2024]. DIETER is a linear cost-minimizing model that abstracts from discrete capacity decisions and economies of scale as well as non-convex operational constraints. The model is centered around the power sector, modeling the target year 2050.
4.1.1 Geographical scope and grid assumptions
The model includes a network of 28 countries (EU27 excluding Malta and Cyprus plus Norway, United Kingdom and Switzerland). Transmission network capacities are on a net transfer (NTC) basis and fixed to the TYNDP 2022 Reference Grid [entso-e_tyndp_2022] (Figure SI.8, left panel).
4.1.2 Generation technologies
The model only permits renewable generation technologies including solar photovoltaics, onshore and offshore wind, hydro generation and some biomass combustion as well as gas turbines running on hydrogen. We only admit nuclear generation capacity in a sensitivity (see Section 2.5). We assume high technical capacity potentials for solar and onshore wind [trondle_home-made_2019], see Table SI.1. Offshore wind capacities are additionally constrained to remain in the range of the Ten-Year Network Development Plan (TYNDP 2024) of the European Network of Transmission System Operators for Electricity (ENTSO-E) [entso-e_tyndp_2025]. Biomass capacities are to remain at most at today’s levels and total annual generation is constrained to that of 2022 [ember_yearly_2024]. Hydro capacities, including reservoirs and run-of-river, are fixed at today’s levels. Today’s capacities are based on the European Resource Adequacy Assessment (ERAA) [de_felice_entso-e_2022].
4.1.3 Storage technologies
Storage technology candidates are limited to lithium-ion batteries and pumped-hydro storage (PHS) for shorter periods. Long-duration storage options are hydrogen tanks, hydrogen caverns, and pit thermal energy storage, described in more detail below. The model chooses battery capacities endogenously, while PHS is fixed to today’s levels [de_felice_entso-e_2022], see Table SI.2.
4.1.4 Demand
The demand side of the power system balance consists of an exogenous and an endogenous part. The exogenous part features a historical base electricity demand time series from Open Power System Data [muehlenpfordt_open_2020] that has been corrected for existing electrified heat demand using data from When2Heat [ruhnau_update_2022]. We add incremental industrial electricity demand resulting from projected process electrification until 2050 [neumann_potential_2023] assuming a flat profile and, for computational reasons, we also assume a flat demand profile to represent electric vehicle demand in 2050 based on the size of today’s car fleet. The implicit assumption is that the short-term variations in mobility demand and the flexibility potential of electric vehicles do not interact substantially with long-duration storage requirements beyond the load they add to the system. The left panel of Figure SI.9 gives the resulting (exogenous) annual electricity demand by country.
The endogenous electricity demand comes from storage technologies, hydrogen electrolysis and heat pumps, either decentrally or from large-scale heat pumps connected to district heating networks. A minor share of electricity demand comes from compressors in the hydrogen network, as explained below.
Apart from an electricity balance, the model also imposes balances for hydrogen and decentral heat as well as district heating networks.
4.1.5 Hydrogen
A country’s hydrogen supply can either be met by domestic production using proton exchange membrane (PEM) electrolysis, imports from other countries within, or imports from countries outside the scope of the model. In the case of the latter, imports can arrive by ship or by pipeline. Such external imports are assumed to be possible only under long-term contracts (LTC) that provide a baseload supply with some hourly flexibility of +/- 10%. We assume that shipped hydrogen comes in as cracked ammonia [spatolisano_ammonia_2023]. Import and interconnection capacities are taken from TYNDP 2024, see Figure SI.8, right panel. We assume LTC import prices of 86.8 €/MWh for shipped hydrogen (ammonia + cracking) and 42-51 €/MWh for pipeline imports [entso-e_tyndp_2025]. To use the network, hydrogen generated by electrolysis needs to be compressed, which creates electricity demand. Longer transport distances requires booster compression, resulting in additional electricity demand.
Hydrogen is stored in either salt caverns underground or more expensive pressurized tanks. For both storage options additional compression is necessary. Cavern storage potentials by country are taken from [caglayan_technical_2020] (see Figure SI.10, right panel). Industrial demand for hydrogen is assumed to have a baseload profile and annual quantities are taken from TYNDP 2024 (see Figure SI.9, middle panel). We assume that all other uses of hydrogen, and especially its derivatives, will be covered by additional imports and do not enter the hydrogen balance. Another source of hydrogen demand stems from hydrogen gas turbines that combust hydrogen to produce electricity. The combination of electrolysis, cavern storage and hydrogen turbines form the long-duration electricity storage (LDES) discussed in this paper.
4.1.6 Heating
The model includes the share of the heating sector that is assumed to be electrified with heat pumps. We assume the use of air-sourced heat pumps. The efficiency of heat pumps varies with the outside temperature, see Figure SI.11, lower panel. For decentralized heat pumps, we set the heat pump capacity exogenously to meet the peak heat demand. We assume that decentralized are equipped with a small buffer storage with a duration of 1.5 hours [roth_power_2024].
[…]
4.2 Data sources
4.2.1 Cost parameters
Most cost parameters are taken from the Danish Energy Agency’s (DEA) forecasts for 2050 [danish_energy_agency_technology_2024]. We provide a detailed list in Table SI.3. All costs and prices are in 2022€. The value of lost load is set at 100,000 €/MWh.
4.2.2 Weather data
We use the largest complete dataset of weather inputs for energy system models that is openly available at the time of writing [antonini_weather-_2024]. We use data from 1940 to 2018. The dataset includes country-level capacity factors for solar photovoltaics, wind onshore and wind offshore as well as inflow data for run-of-river, reservoirs and open pumped-hydro storage. Importantly, it also includes residential and commercial heat demand data for all European countries. The data is based on the ERA-5 reanalysis data of the Copernicus project [hersback_era5_2023] and we use temperature data from the same dataset to generate coefficient of performance time series using the When2Heat tool [ruhnau_update_2022]. We have corrected heat demand and hydro time series for long-run trends [gotske_designing_2024]. We provide details on our approach in SI.XX. Finally, we scale heat demand by annual data from [rozsai_m_jrc-idees-2021_2024]. The resulting annual demands are shown in Figure SI.9, right panel. We note that total demands and implicit heating thresholds reflect today’s building stock. Large-scale insulation and retrofitting could significantly reduce effective useful heating demand in 2050 [zeyen_mitigating_2021].
For the scenarios with central heating networks, we apply the country-wise shares of district heating of total heat demand projected in an optimistic scenario for 2050 by [fallahnejad_district_2024] (Figure SI.10, left panel). For each hour, we apply these shares to the hourly heat demand and assume that the remainder up to 80% is to be supplied by decentral heat pumps, effectively creating two heat demand time series per country.
4.3 Scenario design
We investigate the role of electrified heating for long-duration storage requirements using four main scenarios. Table 1 provides descriptions for each of them.
We assume that 80% of space and water heat demand in residential and commercial buildings (not industry) will be covered with electrified heating options, namely decentralized small-scale heat pumps and large-scale heat pumps within district heating networks. These assumptions conservative in some respects and progressive in others.
The assumed share of 80% is surely at the upper end of range of projected electrification shares. A study by the Joint Research Center (JRC) of the European Commission [tsiropoulos_towards_2020] gives an overview of the final energy consumption in the EU28 buildings sector of different scenarios that reach at least 90% emission reduction by 2050. The share of electricity and ambient ranges around 60% for most scenarios, yet reaches up to 80% for others. The leading German energy transition policy research consortium considers various scenarios for the building sector in Germany, in the majority of which decentralized heat pumps and district heating cover more than 80% of the total heat demand by 2045 [luderer_energiewende_2025].
At the same time, heat electrification only based on heat pumps is likely to be conservative in terms of the demands on the power sector, as it is more efficient than alternatives such as resistive heaters or even hydrogen boilers [rosenow_meta-review_2024].
Data availability
Results data are curated in a Zenodo repository (https://zenodo.org/records/15482638). Technology input data can be reproduced by an auxiliary repository (https://gitlab.com/diw-evu/dieter_public/dieterdata). The weather input data is compiled in the Gitlab repository linked below.
Code availability
The project code is openly available in a Gitlab repository (https://gitlab.com/diw-evu/projects/long-thermal). The analysis is written in Julia and Python and the README file explains the intended order of execution.
Acknowledgments
We acknowledge a research grant from the German Federal Ministry of Education and Research via the Ariadne projects (FKZ 03SFK5NO-2).
CRediT author contributions
Felix Schmidt: Conceptualization, methodology, model development, software, analysis, data curation, visualization, writing — original draft, review and editing
Alexander Roth: Conceptualization, methodology, analysis, writing — original draft, review and editing
Wolf-Peter Schill: Conceptualization, methodology, analysis, writing — review and editing
[…]
Supplemental Information
SI.4 Additional model output
SI.4.1 Generation portfolio
Figure SI.1: Distribution of aggregate renewable generation capacities under different scenarios.
SI.4.2 Spatial dimension
Figure SI.2: Geographical distribution of LDES capacities - No Heat: (A) Distribution of cavern, electrolysis and h2 turbine capacities by country across weather years (1940/41 - 2017/18), (B) - (D) Mean capacities
Figure SI.3: Geographical distribution of LDES capacities - Decent: (A) Distribution of cavern, electrolysis and h2 turbine capacities by country across weather years (1940/41 - 2017/18), (B) - (D) Mean capacities
Figure SI.4: Geographical distribution of LDES capacities - District & Decent: (A) Distribution of cavern, electrolysis and h2 turbine capacities by country across weather years (1940/41 - 2017/18), (B) - (D) Mean capacities
Figure SI.5: Geographical distribution of generation capacities - No Heat: (A) Distribution of solar PV, wind onshore and windoff offshore by country across weather years (1940/41 - 2017/18), (B) - (D) Mean capacities
Figure SI.6: Geographical distribution of generation capacities - Decent: (A) Distribution of solar PV, wind onshore and windoff offshore by country across weather years (1940/41 - 2017/18), (B) - (D) Mean capacities
Figure SI.7: Geographical distribution of generation capacities - District & Decent: (A) Distribution of solar PV, wind onshore and windoff offshore by country across weather years (1940/41 - 2017/18), (B) - (D) Mean capacities
SI.5 Input data
SI.5.1 Networks, demands and potentials
Figure SI.8: Assumed network capacities:
Net transfer capacities in the model are fixed to TYNDP 2022 levels for the electricity grid [entso-e_tyndp_2022] and to TYNDP 2024 levels for the hydrogen network [entso-e_tyndp_2025].
Figure SI.9: Annual energy demands by sector:
(left panel) Exogenous electricity demand consisting of baseline electricity demand based on 2017 historical data from Open Power System Data [muehlenpfordt_open_2020], corrected for existing electrified heat demand using data When2Heat data [ruhnau_update_2022]. ( (Middle panel): Exogenous hydrogen demand from industry based on TYNDP 2024 [entso-e_tyndp_2025]. (Right panel) Mean aggregate heat demands and year-by year variations based on dataset by Antonioni et al. [antonini_weather-_2024] and the Joint Research Council IDEES database [rozsai_m_jrc-idees-2021_2024].
Figure SI.10: District heating and cavern storage potentials:
Left panel: Projected district heating potential by country as share of total heat demand, from [fallahnejad_district_2024]; Right panel: Cavern storage potentials in chemical terawatt hours, from [caglayan_technical_2020].
SI.5.2 Weather data
We use an open-source dataset of 78 weather years with country-level renewable energy capacity factors for solar PV, wind onshore and wind offshore [antonini_weather-_2024]. The dataset also contains hydro inflows for run-of-river dams and basins. Finally, it also provides temperature-dependent space and water heating profiles for European countries.
The dataset does not contain coefficients of performance for heat pumps. The meteorological data underlying the dataset are from the Copernicus Project’s ERA-5 reanalysis database [hersback_era5_2023]. We use the same database to download hourly near-ground temperatures at a granular spatial level. We compute and aggregate coefficients of performance for air-sourced heat pumps to country-level using the When2Heat tool [ruhnau_update_2022] (Figure SI.11, lower panel).
We scale the heat demand profiles by annual country-level heat demands from 2015 as per the JRC-IDEES database [rozsai_m_jrc-idees-2021_2024]. The resulting annual data exhibits a trend, which may reflect both climate change and increased energy efficiency. As we are interested in interannual variations, we estimate a linear model per country to de-trend the data around the 2015 anchor year. The resulting heat demand time series are shown in Figure SI.11, upper panel.
For hydro inflow time series, we following Goetske et al. in correcting inflows around a historical break point [gotske_designing_2024], which we assume to be 1980. Anything prior to that break point will be scaled by the ratio of means of annual inflows of the time series before and after the break point.
Finally, we pro-rate the hydro inflow time series by capacity to distribute inflows between open pumped-hydro plants and pure reservoirs in case both exist in a country.
Figure SI.11: Heat demand and coefficients of performance
SI.6 Technology data
Country Minimum Capacity (GW) Maximum Capacity (GW)
Biomass Nuclear Offshore Wind Onshore Wind Run-of-River Solar PV Biomass Nuclear Offshore Wind Onshore Wind Run-of-River Solar PV
AT 0.0 0.0 0.0 9.0 6.1 12.0 0.6 0.0 0.0 142 6.1 879
BE 0.0 0.0 4.4 4.7 0.2 10.4 0.9 0.0 8.0 38 0.2 499
BG 0.0 2.0 0.0 1.0 0.5 3.2 0.3 2.0 0.6 254 0.5 3484
CZ 0.0 4.1 0.0 1.0 0.4 3.9 0.8 4.1 0.0 203 0.4 2201
DE 0.0 0.0 23.6 75.4 4.7 96.1 7.6 0.0 74.6 684 4.7 8970
DK 0.0 0.0 6.8 6.2 0.0 6.5 2.3 0.0 129.2 42 0.0 1269
EE 0.0 0.0 0.2 0.8 0.0 0.4 0.4 0.0 10.0 180 0.0 967
ES 0.0 3.0 0.2 48.4 3.6 38.4 1.7 3.0 17.4 1032 3.6 18143
FI 0.0 4.4 1.5 12.4 0.0 1.5 3.7 4.4 65.0 1714 0.0 5191
FR 0.0 58.2 5.5 35.9 13.6 43.4 2.4 58.2 60.0 734 13.6 15942
GR 0.0 0.0 0.1 6.9 0.4 7.5 0.5 0.0 22.2 314 0.4 2214
HR 0.0 0.0 0.0 2.8 0.4 0.4 0.4 0.0 3.0 130 0.4 1346
HU 0.0 3.1 0.0 0.3 0.1 6.5 0.9 3.1 0.0 88 0.1 4701
IE 0.0 0.0 3.5 5.8 0.2 0.6 0.1 0.0 35.0 111 0.2 2330
IT 0.0 0.0 0.9 18.4 6.3 51.1 4.7 0.0 35.4 499 6.3 4942
LT 0.0 0.0 0.7 1.5 0.1 0.9 0.2 0.0 7.6 168 0.1 1911
LU 0.0 0.0 0.0 0.4 0.0 0.6 0.1 0.0 0.0 6 0.0 59
LV 0.0 0.0 0.5 0.4 1.6 0.0 0.2 0.0 14.0 251 1.6 1482
NL 0.0 0.5 11.5 8.0 0.0 27.3 3.9 0.5 72.5 25 0.0 722
NO 0.0 0.0 0.2 6.1 0.0 0.6 0.0 0.0 18.5 1105 0.0 8494
PL 0.0 0.0 5.9 8.7 0.5 5.1 1.5 0.0 35.0 732 0.5 9949
PT 0.0 0.0 0.3 8.9 0.8 7.5 1.1 0.0 10.0 142 0.8 2401
RO 0.0 2.0 0.0 5.3 3.4 5.1 0.1 2.0 4.0 465 3.4 8505
SE 0.0 5.9 1.0 16.9 0.0 5.4 4.6 5.9 72.0 2356 0.0 7294
SI 0.0 0.7 0.0 0.2 1.3 1.9 0.1 0.7 0.0 52 1.3 127
SK 0.0 2.7 0.0 0.5 1.6 1.2 0.6 2.7 0.0 136 1.6 1202
CH 0.0 1.2 0.0 0.3 4.2 9.8 0.4 1.2 0.0 43 4.2 178
UK 0.0 9.3 34.8 26.6 2.1 23.4 6.4 9.3 116.9 542 2.1 6612
Table SI.1: Capacity bounds of generation capacities
Country Power in (GW) Power out (GW) Energy (GWh)
Pumped hydro closed Pumped hydro open Pumped hydro closed Pumped hydro open Reservoir Pumped hydro closed Pumped hydro open Reservoir
AT 0.30 5.21 0.30 6.04 2.49 2 1733 757
BE 1.15 1.22 5
BG 0.78 0.15 0.86 0.54 1.28 9 255 843
CH 1.90 2.05 1.90 10.73 70 8800
CZ 0.66 0.44 0.69 0.47 0.70 4 2 3
DE 7.42 1.36 7.38 1.64 1.30 242 417 259
DK
EE
ES 6.68 2.42 6.87 2.68 10.97 99 5962 11649
FI 3.24 5530
FR 1.95 1.85 1.95 1.85 8.79 10 90 10000
GR 0.71 0.74 0.68 0.70 2.74 5 3600
HR 0.25 0.28 1.94 18 2046
HU
IE 0.29 0.29 2
IT 7.37 2.10 7.30 3.33 9.65 70 309 5649
LT 0.90 0.93 11
LU 1.03 1.31 5
LV
NL
NO 1.09 38.05 89577
PL 1.49 0.17 1.32 0.22 0.27 6 2 1
PT 3.58 3.83 3.77 2018 1290
RO 0.10 0.84 2.58 1506 2425
SE 16.45 31931
SI 0.19 0.18 3
SK 0.65 0.18 0.65 0.28 0.02 4 46 17
UK 2.68 2.95 26
Table SI.2: Capacities of hydro plants
Technology Costs Efficiency Lifetime Electricity demand
Investment Fixed O&M Variable+Fuel
[€/kW or €/kWh] [€/kW/a or €/kWh/a] [€/MWh] [%] [years] [of input]
Biomass 2517 129.01 13.65 48.7% 30
Nuclear 8080 102.62 5.24 33.7% 40
H2 turbine 501 7.893 $p_{H2}$ + 5.00 43.0% 25
Offshore wind 1843 37.08 3.37 30
Onshore wind 1225 17.53 2.08 30
Run-of-river 3870 77.41 50
Solar PV 458 8.76 40
Battery inverter 72 0.65 2.00 96.0% 30
Battery storage 112 99.9% 30
PEM electrolysis 404 8.05 66.4% 20/25 0.75%
Hydrogen cavern storage (kWhh2) 1.43 100% 100
H2 cavern compressor (kWh2) 96 3.83 15 2.10%
Hydrogen tank storage (kWhh2) 17 0.48 30
H2 tank compressor (kWh2) 10 0.48 25 6.00%
Large-scale air-sourced heat pump (kWth) 970 2.55 COP 25
Pit thermal storage (kWhth) 0.57 0.0037 49.12%* 25
Table SI.3: Technology cost assumptions
The data is taken from the Technology Catalogues of the Danish Energy Agency [danish_energy_agency_technology_2024]. For pit thermal storage efficiency, the number reflects the efficiency for a 90 day storage duration accounting for heat loss and (dis-)charge losses.

### v2

@@
 Yet, most studies at the nexus of LDES and weather variability focus only on the power sector and neglect the impact of weather variations on the demand side of a sector-coupled renewable system [gotske_designing_2024].
-Accounting for 16% percent of total greenhouse gas emissions in the European Union in 2024 [rozsai_m_jrc-idees-2021_2024], the heating sector takes center stage in the continent’s decarbonization efforts [commission_renewable_2023].
+Accounting for 16% of total greenhouse gas emissions in the European Union in 2024 [rozsai_m_jrc-idees-2021_2024], the heating sector takes center stage in the continent’s decarbonization efforts [commission_renewable_2023].
 Previous studies have shown that serving a substantial part of residential and commercial heat demand with heat pumps can have challenging consequences for the power system [bloess_power--heat_2018, zeyen_mitigating_2021, roth_power_2023, roth_power_2024].
@@
 2 Results
-In the following, we focus primarily on the link between electrification of residential and commercial heat and LDES needs.
+In the following, we focus primarily on the link between the electrification of residential and commercial heat and LDES needs.
 The Supplemental Information Section SI.4 provides several additional results.
@@
 In 1962/63, the year with the highest LDES requirement in the set of runs with electrified heat, the LDES capacity increases tenfold from 40 TWh to 400 TWh.
-In Figure 1, the weather years are ordered by LDES requirement in the Decent scenario.
-Notably, the No Heat scenario does not reflect the same ordering.
+In Figure 1, the weather years are sorted by LDES requirement in the Decent scenario.
+Notably, the No Heat scenario does not reflect the same order.
 Weather years that have high LDES requirements in the scenario without electrified heat do not necessarily rank high when considering electrified heat (cf.
@@
 Optimal long-duration electricity storage capacities by weather year.
-Grey dots represent optimal capacities in the No Heat scenario.
+Gray dots represent optimal capacities in the No Heat scenario.
 Dark-red dots represent optimal capacities in the Decent scenario.
@@
 The results of these model runs (light red crosses in Figure 1) highlight the drivers of increasing LDES requirements.
-Heat demand is seasonal and electrified heating naturally adds most load in winter months.
+Heat demand is seasonal, and electrified heating naturally adds most load in winter months.
 As extreme system states of VRE scarcity mostly occur during winter months [kittel_quantifying_2024, kittel_coping_2025, gotske_designing_2024], the additional load amplifies the energy deficit, triggering additional LDES requirements by functioning as a lever.
 Across the sample of weather years, this leverage effect accounts for 98 TWh or 75% of the total LDES increase on average.
-For some years, the Decent - Mean scenario results in even higher LDES requirements (e.g., 1973/74) in which unusually mild conditions offset some of this leverage effect.
-Yet, in other years the year-specific heat demand leads to far higher LDES capacities, as VRE scarcities coincide with cold spells in a compound effect.
+For some years, the Decent - Mean scenario results in even higher LDES requirements (e.g., 1973/74) for which unusually mild conditions offset some of this leverage effect.
+Yet, in other years, the year-specific heat demand leads to far higher LDES capacities, as VRE scarcities coincide with cold spells in a compound effect.
 While this effect only explains 25% on average, cold spells have a considerable impact in some of the most extreme weather years.
@@
 LDES trajectories shown against daily deviations in heat demand from long-run mean (bright red - mid) and 24-hour rolling renewable availability (green - low).
-Gray area represents storage trajectory in No Heat.
+The gray area represents the storage trajectory in No Heat.
 The darker red area shows the additional storage level in Decent - Mean, and the lighter red addition represents the incremental storage level from the specific weather year’s heat demand (Decent).
 We illustrate both effects for four weather years with high additional LDES requirements in Figure 2.
-In the deterministic setting addressed here, the maximum required storage level over the course of a weather year determines the optimal LDES energy capacity.
-The trajectories therefore provide an insight into the critical periods that determine optimal LDES capacities.
+In the deterministic setting addressed here, the maximum required storage level in the course of a weather year determines the optimal LDES energy capacity.
+The trajectories, therefore, provide an insight into the critical periods that determine optimal LDES capacities.
 The gray areas correspond to the storage levels observed in the scenario without electrified heat (No Heat).
@@
 The lighter red areas show the storage trajectories arising from the weather year’s specific heat demand profile (Decent).
-In the lower part of the figure, we plot the absolute daily deviations from long-run mean profiles of both heat demand (in dark-red) and renewable generation.
+In the lower part of the figure, we plot the absolute daily deviations from long-run mean profiles of both heat demand (in dark red) and renewable generation.
 The heat deviations serve as a proxy for cold spells.
@@
 The weather year of 1958/59 (center right) results in the highest LDES capacity in the No Heat scenario due to a severe episode of energy scarcity between November and early January, leading to a discharge event of 79 TWh over the course of 28 days.
-Rather mild temperatures induce a LDES increase almost entirely through the leverage effect, where the discharge event increases in size to 223 TWh end extends to 106 days.
+Rather mild temperatures induce a LDES increase almost entirely through the leverage effect, where the discharge event increases in size to 223 TWh and extends to 106 days.
 Lastly, the season of 1941/42 (right) exhibits one of the lowest LDES requirements across all years in the sample when we do not include electrified heat.
@@
 […]
-In contrast to a hydrogen cavern storage, a LDTS has relatively high standing losses from heat loss [sifnaios_heat_2023].
+In contrast to a hydrogen cavern storage, an LDTS has relatively high standing losses from heat loss [sifnaios_heat_2023].
 In the base specification, we assume an efficiency equivalent to cumulative losses of 39% within 90 days of storage in energetic terms [zeyen_mitigating_2021].
@@
 The introduction of LDTS reduces the additional LDES requirements from electrified heat by 60 TWh (36%) on average (Figure 3, left panel).
-In the extreme weather year 1962/63 the introduction of co-optimized LDTS capacities in DH networks leads to a reduction of 155 TWh (37%).
+In the extreme weather year 1962/63, the introduction of co-optimized LDTS capacities in DH networks leads to a reduction of 155 TWh (37%).
 The Europe-wide average LDTS capacity over all years is 143 TWh, reaching almost 323 TWh in the maximum (Figure 3, center panel).
@@
 Any of those countries could technically host all the cavern storage required for the entire continent.
-Yet, in the Decent scenario around 43% (72 out of 168 TWh) of the total cavern storage capacity is located in Germany on average, compare Figure SI.3.
+Yet, in the Decent scenario, around 43% (72 out of 168 TWh) of the total cavern storage capacity is located in Germany on average, compare Figure SI.3.
 Notably, cavern storage is much more evenly distributed between countries in the No Heat scenario (Figure SI.2, left panel).
@@
 Figure 4 shows the change in the geographical long-duration storage distribution across Europe when district heating networks with LDTS are introduced for the example of 1962/63.
-Heating demands and district heating potentials vary between countries such that a meaningful comparison of LDTS deployment requires a relative basis.
+Heating demands and district heating potentials vary between countries, such that a meaningful comparison of LDTS deployment requires a relative basis.
 We express LDTS as a share of total demand in district heating networks.
 LDTS uptake is particularly high in Central Eastern Europe where there is very little LDES capacity.
-The reduction in LDES capacity is disproportionally high in Germany.
+The reduction in LDES capacity is disproportionately high in Germany.
 Consequently, the share of European LDES capacity in Germany falls to 35% (38 out of 108 TWh) in District & Decent.
@@
 For one, Germany has a relatively high district heating potential compared to other cavern storage locations, which means that it can substitute some of its on flexibility requirements locally.
-For two, German LDES provides flexibility in other countries, particularly in Central Eastern Europe and high LDTS adoption in these countries also reduces German LDES requirements.
+For another, German LDES provides flexibility in other countries, particularly in Central Eastern Europe and high LDTS adoption in these countries also reduces German LDES requirements.
 Figure 4: Geographical impact of adding long-duration thermal storage in 1962/63
@@
 For each of the four weather years, we run 42 alternative settings, varying the LDTS standing losses on the x-axis and the investment costs on the y-axis (Figure 5).
-The base configuration used in all prior scenarios above is in the corner in the bottom-left.
+The base configuration used in all prior scenarios above is in the bottom-left corner.
 In the most favorable parameter configurations LDTS reduces LDES needs by another 19-35% (upper panel).
-In the case of 1941/42 this would correspond to an LDES requirement of 111 TWh when LDTS is free and efficiency is at 99%.
+In the case of 1941/42, this would correspond to an LDES requirement of 111 TWh when LDTS is free and efficiency is at 99%.
 As per the lower panel of Figure 5, such additional reductions are accompanied by large increases in optimal LDTS capacities.
 For 1941/42, the LDTS capacity almost triples between the base case and the best parameter combination.
-Based on the conclusion above that even in the base specification heat losses do not match the roundtrip losses of LDES for relevant storage durations, it is not surprising that additional substitution from efficiency gains beyond a threshold is limited.
-In the top line of each tile LDTS can be built for free.
+Based on the conclusion above that even in the base specification, heat losses do not match the round-trip losses of LDES for relevant storage durations, it is not surprising that additional substitution from efficiency gains beyond a threshold is limited.
+In the top line of each tile, LDTS can be built for free.
 In these cases, the substitution potential with respect to LDES is limited by the country-level DH potentials.
@@
 The x-axis varies heat losses from the base case to almost no losses in energetic terms, expressed as energy remaining after 90 days in storage.
-The y-axis varies the assumed investment costs of LDTS by multiplying them with the labeled scalers.
+The y-axis varies the assumed investment costs of LDTS by multiplying them by the labeled scalers.
 Next, we explore an extreme counterfactual case in which all electrified heating would be supplied through heating networks (Table 2, last row), which makes LDTS available for all electrified heating demand.
@@
 In the District & Decent base case, hydrogen imports with long-term contracts are actually slightly higher than in the Decent base case, which is likely to replace some LDES capacity otherwise used for serving flat industrial hydrogen demand.
-In absence of hydrogen import capacities, this option no longer exists.
+In the absence of hydrogen import capacities, this option no longer exists.
 In the scenario with neither hydrogen nor electricity interconnection capacity (’Islands’, rows 4 and 8), LDES capacities increase only marginally compared to the ’No H2 network’ case and even decrease in the case of 1996/97 in the Decent scenario, indicating that the system value of LDES depends on interconnection.
 Without interconnection, building larger generation portfolios is cheaper than building expensive hydrogen tank storage.
-Only about 10 TWh of tank storage are being built across Europe in the ’Islands’ case of the Decent scenario.
+Only about 10 TWh of tank storage is being built across Europe in the ’Islands’ case of the Decent scenario.
 Lastly, we test the effect of introducing around 97 GW of nuclear capacity, of which the majority is located in France (58 GW).
@@
 The reason is likely to be twofold.
-For one, nearly two thirds of the capacity being in France, limited interconnection capacities may prevent spatial balancing during extreme system episodes.
+For one, nearly two-thirds of the capacity is in France, and limited interconnection capacities may prevent spatial balancing during extreme system episodes.
 By contrast, LDES leaves the option to transport both electricity and hydrogen.
@@
 We further use a simplified representation of the thermodynamics of large-scale pit thermal storage, which neglects that some electricity might be required at the time of LDTS discharging to elevate the temperature level again, which could reduce the cost advantage of thermal storage against hydrogen storage.
-While some of these assumptions (e.g., only air-sourced heat pumps, no other heat sources) lead to a higher electricity demand and higher LDES needs than in reality expected, others might lead to an underestimation (simplified representation of the thermodynamics).
+While some of these assumptions (e.g., only air-sourced heat pumps, no other heat sources) lead to a higher electricity demand and higher LDES needs than expected in reality, others might lead to an underestimation (simplified representation of the thermodynamics).
 Finally, the perfect foresight assumption for a given weather year leads to ideal but unrealistic storage operations and is likely to underestimate the relative value of LDES.
@@
 Presently, bar a few test sites, hydrogen cavern storage is a nascent technology.
-Long permitting processes and underdeveloped supply chains, as well as long-term uncertainty about hydrogen prices and quantities are likely to increase project risks [steinbach_future_2024].
+Long permitting processes and underdeveloped supply chains, as well as long-term uncertainty about hydrogen prices and quantities, are likely to increase project risks [steinbach_future_2024].
 Similarly, neither electrolysis nor hydrogen gas turbines have been deployed at scale so far [odenweller_green_2025].
@@
 In this context, LDTS could provide a promising complementary technology, as it does not rely on particular geological conditions and can effectively deal with long-duration flexibility requirements in the heating sector.
-Yet, it is likely to be confined to heating networks and our results suggest that even if they were substantially expanded, a sizable residual long-duration flexibility need from the heating sector would remain.
+Yet, it is likely to be confined to heating networks, and our results suggest that even if they were substantially expanded, a sizable residual long-duration flexibility need from the heating sector would remain.
 Our results indicate that LDES in the form of cavern storage would be concentrated in a few countries, particularly in Germany, and thus other countries without salt cavern potentials to store hydrogen would be dependent on imports and transmission infrastructure to cover their load and heat demand.
-As this could raise regulatory and geopolitical questions, LDTS could for some countries increase their energy autonomy and decrease dependence on energy transport infrastructure.
+As this could raise regulatory and geopolitical questions, LDTS could, for some countries, increase their energy autonomy and decrease dependence on energy transport infrastructure.
 Accordingly, LDES and LDTS should not be considered as substitutes, but rather complement each other, and policymakers should work to enable the mass deployment of both storage types.
 In particular, policymakers and regulators should expedite the process of designing adequate regulatory frameworks that enable a timely capacity build-out, as a lack of long-duration flexibility would jeopardize the electrification of the heating sector.
-From an investor’s perspective, the increased variance of optimal LDES capacities between weather years when considering electrified heating raises additional questions around predictable cost recovery without other economic incentives.
+From an investor’s perspective, the increased variance of optimal LDES capacities between weather years, when considering electrified heating, raises additional questions around predictable cost recovery without other economic incentives.
 This clearly calls for de-risking such investments.
@@
 Finally, the use of deterministic models omits a potentially important advantage of LDES over LDTS in settings without perfect foresight.
-LDES can serve final energy demand in many different sectors depending on the realized system state thanks to the usefulness and convertibility of hydrogen.
+LDES can serve final energy demand in many different sectors, depending on the realized system state, thanks to the usefulness and convertibility of hydrogen.
 In contrast, LDTS is less useful to the energy system as it can only serve heat demand.
-Future research, using stochastic modelling (such as in [schmidt_long-duration_2025]) could be a starting point to better understand the value of different types of storage, the interaction between them, and their operation in a non-deterministic setting.
+Future research, using stochastic modeling (such as in Schmidt [schmidt_long-duration_2025]) could be a starting point to better understand the value of different types of storage, the interaction between them, and their operation in a non-deterministic setting.
 However, such approaches are likely to be feasible only for less extensive geographical and temporal scopes than modeled here.
@@
 4.1.1 Geographical scope and grid assumptions
-The model includes a network of 28 countries (EU27 excluding Malta and Cyprus plus Norway, United Kingdom and Switzerland).
+The model includes a network of 28 countries (EU27 excluding Malta and Cyprus plus Norway, the United Kingdom and Switzerland).
 Transmission network capacities are on a net transfer (NTC) basis and fixed to the TYNDP 2022 Reference Grid [entso-e_tyndp_2022] (Figure SI.8, left panel).
@@
 The exogenous part features a historical base electricity demand time series from Open Power System Data [muehlenpfordt_open_2020] that has been corrected for existing electrified heat demand using data from When2Heat [ruhnau_update_2022].
-We add incremental industrial electricity demand resulting from projected process electrification until 2050 [neumann_potential_2023] assuming a flat profile and, for computational reasons, we also assume a flat demand profile to represent electric vehicle demand in 2050 based on the size of today’s car fleet.
+We add incremental industrial electricity demand resulting from projected process electrification until 2050 [neumann_potential_2023], assuming a flat profile and, for computational reasons, we also assume a flat demand profile to represent electric vehicle demand in 2050 based on the size of today’s car fleet.
 The implicit assumption is that the short-term variations in mobility demand and the flexibility potential of electric vehicles do not interact substantially with long-duration storage requirements beyond the load they add to the system.
@@
 To use the network, hydrogen generated by electrolysis needs to be compressed, which creates electricity demand.
-Longer transport distances requires booster compression, resulting in additional electricity demand.
+Longer transport distances require booster compression, resulting in additional electricity demand.
 Hydrogen is stored in either salt caverns underground or more expensive pressurized tanks.
-For both storage options additional compression is necessary.
+For both storage options, additional compression is necessary.
 Cavern storage potentials by country are taken from [caglayan_technical_2020] (see Figure SI.10, right panel).
-Industrial demand for hydrogen is assumed to have a baseload profile and annual quantities are taken from TYNDP 2024 (see Figure SI.9, middle panel).
+Industrial demand for hydrogen is assumed to have a baseload profile, and annual quantities are taken from TYNDP 2024 (see Figure SI.9, middle panel).
 We assume that all other uses of hydrogen, and especially its derivatives, will be covered by additional imports and do not enter the hydrogen balance.
 Another source of hydrogen demand stems from hydrogen gas turbines that combust hydrogen to produce electricity.
-The combination of electrolysis, cavern storage and hydrogen turbines form the long-duration electricity storage (LDES) discussed in this paper.
+The combination of electrolysis, cavern storage and hydrogen turbines forms the long-duration electricity storage (LDES) discussed in this paper.
 4.1.6 Heating
@@
 For decentralized heat pumps, we set the heat pump capacity exogenously to meet the peak heat demand.
-We assume that decentralized are equipped with a small buffer storage with a duration of 1.5 hours [roth_power_2024].
+We assume that decentralized heat pumps are equipped with a small buffer storage with a duration of 1.5 hours [roth_power_2024].
 […]
@@
 Importantly, it also includes residential and commercial heat demand data for all European countries.
-The data is based on the ERA-5 reanalysis data of the Copernicus project [hersback_era5_2023] and we use temperature data from the same dataset to generate coefficient of performance time series using the When2Heat tool [ruhnau_update_2022].
+The data is based on the ERA-5 reanalysis data of the Copernicus project [hersback_era5_2023] and we use temperature data from the same dataset to generate the coefficient of performance time series using the When2Heat tool [ruhnau_update_2022].
 We have corrected heat demand and hydro time series for long-run trends [gotske_designing_2024].
@@
 For the scenarios with central heating networks, we apply the country-wise shares of district heating of total heat demand projected in an optimistic scenario for 2050 by [fallahnejad_district_2024] (Figure SI.10, left panel).
-For each hour, we apply these shares to the hourly heat demand and assume that the remainder up to 80% is to be supplied by decentral heat pumps, effectively creating two heat demand time series per country.
+For each hour, we apply these shares to the hourly heat demand and assume that the remainder up to 80% is to be supplied by decentralized heat pumps, effectively creating two heat demand time series per country.
 4.3 Scenario design
@@
 We assume that 80% of space and water heat demand in residential and commercial buildings (not industry) will be covered with electrified heating options, namely decentralized small-scale heat pumps and large-scale heat pumps within district heating networks.
-These assumptions conservative in some respects and progressive in others.
-The assumed share of 80% is surely at the upper end of range of projected electrification shares.
+These assumptions are conservative in some respects and progressive in others.
+The assumed share of 80% is surely at the upper end of the range of projected electrification shares.
 A study by the Joint Research Center (JRC) of the European Commission [tsiropoulos_towards_2020] gives an overview of the final energy consumption in the EU28 buildings sector of different scenarios that reach at least 90% emission reduction by 2050.